# Universality in driven systems with a multiply-degenerate umbilic point

### v1

@@ Abstract
We investigate a driven particle system, a multilane asymmetric exclusion process, where the particle number in every lane is conserved, and stationary state is fully uncorrelated. The phase space has, starting from three lanes and more, an umbilic manifold where characteristic velocities of all the modes but one coincide, thus allowing us to study a weakly hyperbolic system with arbitrarily large degeneracy. We then study space-time fluctuations in the steady state, at the umbilic manifold, which are expected to exhibit universal scaling features. We formulate an effective mode-coupling theory (MCT) for the multilane model within the umbilic subspace and test its predictions. Unlike in the bidirectional two-lane model with an umbilic point studied earlier, here we find a robust $z=3/2$ dynamical exponent for the umbilic mode. The umbilic scaling function, obtained from Monte-Carlo simulations, for the simplest 3-lane scenario, appears to have an universal shape for a range of interaction parameters. Remarkably, the shape and dynamic exponent of the non-degenerate mode can be analytically predicted on the base of effective MCT, up to non-universal scaling factor. Our findings suggest the existence of novel universality classes with dynamical exponent $3/2$, appearing in long-lived hydrodynamic modes with equal characteristic velocities.
@@ 1 Introduction
A cornerstone assumption of nonlinear fluctuating hydrodynamics and mode coupling theory is a spatial separation of modes with time, an assumption refereed to as strong hyperbolicity in the theory of hyperbolic conservation laws and hyperbolic shocks [7]. Strong hyperbolicity is crucial for an analysis of hyperbolic shocks, which are, in turn, crucial for understanding intrinsic features of non-equilibrium systems, with no equilibrium counterpart, e.g. boundary-driven phase transitions [4].
@@
Figure 1: Allowable hops for the three-lane model ($K+1=3$) with their rates (2.1). Particles with the same color hop with the same rates, e.g. hopping of black-coloured particles is forbidden by hardcore exclusion.
@@ 2 Multilane TASEP: an ideal candidate for studying umbilic modes
Here we need to stress that (2.30) and (2.33) are effective mode-coupling matrices and respectively (2.19) are effective mode-coupling equations, based on symmetries (2.10) occurring at the umbilic manifold. In the effective MCT (2.19),(2.30) the umbilic mode appears as a non-degenerate mode with self-coupling proportional to $g_{1}$. We emphasize that we do expect (2.19) to describe the umbilic mode qualitatively; for the latter purpose we need full system dynamics, captured by e.g. full Monte-Carlo simulations.
@@
for the dynamical exponent of the umbilic mode. For the remaining single mode, accounting for $\bar{G}^{s}_{11}\sim g_{1}>0$, two scenarios are possible [12]: (A) if $\bar{G}^{s}_{22}\neq 0$ then the single mode is a KPZ mode $z_{s}=3/2$ and $S_{s}\equiv f_{\rm KPZ}$ [2]. (B) if $\bar{G}^{s}_{22}\sim g_{3}=0$, i.e.
@@
which is a maximally asymmetric Levy stable $\frac{5}{3}$ distribution. The only free parameter in (2.39) is a nonuniversal scaling factor $E_{s}$, which cannot be predicted from mode-coupling theory [8, 10].
@@ 3 Simulation results
For the umbilic mode with double degeneracy $K=2$ we find an evidence of the dynamic exponent $z_{u}=3/2$ (2.36); however the umbilic scaling function $f_{u}$ appears very different from the Prahofer-Spohn (KPZ universality) scaling function. Moreover, $f_{u}$ differs from any known scaling function, which can be attributed to $z=3/2$ and is symmetric: not only KPZ, but also e.g. Moore-Colaiori function [14], or symmetric Levy stable $3/2$ distribution. From our numerical data, we can tabulate $f_{u}$ for a doubly degenerate umbilic mode $K=2$ with relatively good precision. Remarkably, the shape of the umbilic function is robust to changes of system parameters, i.e. appears universal on the whole umbilic line, see Fig. 6. A possible reason for this universality may be the following: curiously, our umbilic mode considered alone, i.e. ignoring the nondegenerate mode, falls into the umbilic scenario of $2$-component systems described in Ref. [17], and can be classified according to it as a special point on a fixed points line, see C for details. Further confirmation of this reasoning is that, within our numerical precision, the scaling function from continuous time Burgers equation analysis in Ref. [17] match our umbilic scaling function determined from Monte-Carlo simulations, see Fig 4.
As far as the nondegenerate mode is concerned, our findings agree well with mode-coupling theory predictions (2.38), (2.39), see Fig. 5. Note that due to absence of the off-diagonal terms in $\bar{G}^{s}$ from (2.33) finite-time effects are small and the Levy stable $5/3$ distribution (2.39) is established already for relatively small times. Within our numerical accuracy, starting from times of order $t=10^{4}$ Monte Carlo steps we see no deviations from the analytical result (2.39) apart from a small deviation in the finite-time dynamical exponent. Our findings for the single mode do not depend on the value of interaction $a$, as long as the (B) condition $g_{3}=0$ (2.37) is satisfied.
Figure 4: Data collapse for the umbilic mode from lattice Monte Carlo simulations with $z_{u}=3/2$, $c_{u}=5/8$ and $E_{u}=2.12$. Shaded regions show two standard deviation neighborhoods. Black curves show the best-fit KPZ (Prähofer-Spohn) scaling function (dashed curve) and the (scaled) $S_{11}$ correlator of Ref. [17] (full curve) at X=Y=1X=Y=-1. Simulation parameters: $K=2$, $a=2$, $\rho=1/4$, system size $N=10^{6}$, $n_{\textrm{samples}}=10^{3}$.
@@ 4 Numerical integration of continuous NLFH equations
Figure 7: Data collapse from continuous NLFH simulation for: (left panel) the umbilic mode with $\overline{z}_{u}=3/2$, $\overline{c}_{u}=0.575$ and $\overline{E}_{u}=2.32$. Black curve shows umbilic scaling function estimated from Monte Carlo simulation for comparison. (right panel) the for heat mode where the best data collapse is obtained for $\overline{z}_{u}=1.74$, $\overline{c}_{u}=1.75$ and $\overline{E}_{u}=0.94$. Black curves show maximally asymmetric Levy stable $\frac{5}{3}$ (full line) and $\overline{z}_{s}$ (dashed line) distributions. Simulation parameters: $K=2$, $\rho=1/4$, $a=2$ with effective mode-coupling matrices $\bar{G}^{u}=\frac{3}{8}\left(\begin{matrix}2&1\\ 1&0\end{matrix}\right)$ and $\bar{G}^{s}=\frac{3}{8}\left(\begin{matrix}2\sqrt{2}&0\\ 0&0\end{matrix}\right)$ and discretization parameters $\Delta t=2\times 10^{-3}$, $L=2^{20}$.
@@
Up to now we only described our numerical results for the double-degenerate $K=2$ umbilic mode, while our effective MCT theory predictions, namely Eqs.(2.36), (2.38), and (2.39) are expected to be valid for an umbilic mode of arbitrary degeneracy $K$. We also did a similar analysis for $K>2$ which is more challenging numerically, when it comes to Monte-Carlo simulations. While the $z_{u}=3/2$ prediction of MCT (2.36) remains robust, the umbilic scaling function $f_{u}$ changes significantly with the umbilic mode degeneracy $K$. With increasing $K$ the umbilic scaling function $f_{u}(K)$ starts closely approaching the KPZ scaling function, data not shown. It would be interesting to investigate if the convergence $f_{u}\rightarrow f_{\rm KPZ}$ takes place in the “meanfield” limit $K\rightarrow\infty$.

### v2

@@ Abstract
We investigate a driven particle system, a multilane asymmetric exclusion process, where the particle number in every lane is conserved, and stationary state is fully uncorrelated. The phase space has, starting from three lanes and more, an umbilic manifold where characteristic velocities of all the modes but one coincide, thus allowing us to study a weakly hyperbolic system with arbitrarily large degeneracy. We then study space-time fluctuations in the steady state, at the umbilic manifold, which are expected to exhibit universal scaling features. We formulate an effective mode-coupling theory (MCT) for the multilane model within the umbilic subspace and test its predictions. Unlike in the bidirectional two-lane model with an umbilic point studied earlier, here we find a robust $z=3/2$ dynamical exponent for the umbilic mode. The umbilic scaling function, obtained from Monte-Carlo simulations appears to have a universal shape for a range of interaction parameters and depends only on umbilic mode degeneracy. Remarkably, the shape and dynamic exponent of the non-degenerate mode can be analytically predicted on the base of effective MCT, up to a non-universal scaling factor. Our findings suggest the existence of novel universality classes with dynamical exponent $3/2$, appearing in long-lived hydrodynamic modes with equal characteristic velocities.
@@ 1 Introduction
A cornerstone assumption of nonlinear fluctuating hydrodynamics and mode coupling theory is a spatial separation of modes with time, an assumption referred to as strong hyperbolicity in the theory of hyperbolic conservation laws and hyperbolic shocks [7]. Strong hyperbolicity is crucial for an analysis of hyperbolic shocks, which are, in turn, crucial for understanding intrinsic features of non-equilibrium systems, with no equilibrium counterpart, e.g. boundary-driven phase transitions [4].
@@
Figure 1: Allowable hops for the three-lane model ($K+1=3$) with their rates (2.1). Particles with the same color hop with the same rates, e.g. hopping of black-coloured particles is forbidden by hardcore exclusion. Note that all particles are identical, and they do not change lanes, so that the overall particle number in each lane is conserved.
@@ 2 Multilane TASEP: an ideal candidate for studying umbilic modes
Here we need to stress that (2.30) and (2.33) are effective mode-coupling matrices and respectively (2.19) are effective mode-coupling equations, based on symmetries (2.10) occurring at the umbilic manifold. In the effective MCT (2.19),(2.30) the umbilic mode appears as a non-degenerate mode with self-coupling proportional to $g_{1}$. We emphasize that we cannot expect (2.19) to describe the umbilic mode quantitatively; for the latter purpose we need full system dynamics, captured by e.g. full Monte-Carlo simulations.
@@
for the dynamical exponent of the umbilic mode. For the remaining single mode, accounting for $\bar{G}^{s}_{11}\sim g_{1}>0$, two scenarios are possible [12]: (A) if $\bar{G}^{s}_{22}\neq 0$ then the single mode is a KPZ mode $z_{s}=3/2$ and $S_{s}\equiv f_{\rm KPZ}$  [^1] Note that the mode-coupling prediction for the scaling function of a single-mode with self-coupling [14] is distinct from $f_{\rm KPZ}$ but is known to be incorrect [2], with the latter being the correct scaling function in this scenario.. (B) if $\bar{G}^{s}_{22}\sim g_{3}=0$, i.e.
@@
which is a maximally asymmetric Levy stable $\frac{5}{3}$ distribution. The only free parameter in (2.39) is a nonuniversal scaling factor $E_{s}$, which cannot be predicted from mode-coupling theory [8, 10]. For the umbilic mode we determine the scaling factor as [2]
$E_{u}=2^{3/2}|G^{u}_{uu}|$ (2.40)
while for the heat mode we determine the scaling factor $E_{s}$ by matching the maxima of the scaling function (2.39).
@@ 3 Simulation results
For the umbilic mode with double degeneracy $K=2$ we find an evidence of the dynamic exponent $z_{u}=3/2$ (2.36); however the umbilic scaling function $f_{u}$ appears very different from the Prahofer-Spohn (KPZ universality) scaling function. Moreover, $f_{u}$ differs from any known scaling function, which can be attributed to $z=3/2$ and is symmetric: not only KPZ, but also the e.g. Moore-Colaiori function [14], or symmetric Levy stable $3/2$ distribution. From our numerical data, we can tabulate $f_{u}$ for a doubly degenerate umbilic mode $K=2$ with relatively good precision. Remarkably, the shape of the umbilic function is robust to changes of system parameters, i.e. appears universal on the whole umbilic line, see Fig. 6. A possible reason for this universality may be the following: curiously, our umbilic mode considered alone, i.e. ignoring the nondegenerate mode, falls into the umbilic scenario of $2$-component systems described in Ref. [17], and can be classified according to it as a special point on a fixed points line, see C for details. Further confirmation of this reasoning is that, within our numerical precision, the scaling function from continuous time Burgers equation analysis in Ref. [17] match our umbilic scaling function determined from Monte-Carlo simulations, see Fig 4.
As far as the nondegenerate mode is concerned, our findings agree well with mode-coupling theory predictions (2.38), (2.39), see Fig. 5. Note that due to absence of the off-diagonal terms in $\bar{G}^{s}$ from (2.33) finite-time effects are small and the Levy stable $5/3$ distribution (2.39) is established already for relatively small times. Within our numerical accuracy, starting from times of order $t=10^{4}$ Monte Carlo steps we see no deviations from the analytical result (2.39) apart from a small deviation in the finite-time dynamical exponent.
Our findings for the single mode do not depend on the value of interaction $a$, as long as the (B) condition $g_{3}=0$ (2.37) is satisfied.
Figure 4: Data collapse for the doubly degenerate (K=2) umbilic mode from lattice Monte Carlo simulations with $z_{u}=3/2$, $c_{u}=5/8$. Shaded regions show two standard deviation neighborhoods. Full black curve, overlapping with our data, shows the (scaled) $S_{11}$ correlator of Ref. [17] at X=Y=1X=Y=-1. Black dashed curve shows the best-fit KPZ (Prähofer-Spohn) scaling function. Simulation parameters: $K=2$, $a=2$, $\rho=1/4$, system size $N=10^{6}$, $n_{\textrm{samples}}=10^{3}$.
@@ 4 Numerical integration of continuous NLFH equations
Figure 7: Data collapse from continuous NLFH simulation for: (left panel) the umbilic mode with $\overline{z}_{u}=3/2$, $\overline{c}_{u}=0.575$ and $\overline{E}_{u}=2.32$ with $\overline{E}_{u}$ determined by matching the height of the umbilic scaling function estimated from Monte Carlo simulations (black curve). (right panel) the for heat mode where the best data collapse is obtained for $\overline{z}_{u}=1.74$, $\overline{c}_{u}=1.75$ and $\overline{E}_{s}=0.94$. Black curves show maximally asymmetric Levy stable $\frac{5}{3}$ (full line) and $\overline{z}_{s}$ (dashed line) distributions. Simulation parameters: $K=2$, $\rho=1/4$, $a=2$ with effective mode-coupling matrices $\bar{G}^{u}=\frac{3}{8}\left(\begin{matrix}2&1\\ 1&0\end{matrix}\right)$ and $\bar{G}^{s}=\frac{3}{8}\left(\begin{matrix}2\sqrt{2}&0\\ 0&0\end{matrix}\right)$ and discretization parameters $\Delta t=2\times 10^{-3}$, $L=2^{20}$.
@@
Figure 8: Data collapse for systems with high (triple $K=3$ (dashed lines) and quadruple $K=4$ (dotted lines)) degeneracy, from lattice Monte Carlo simulations with 44 lanes and 55 lanes respectively. Collapse for the umbilic mode (left panel) with the best-fit KPZ (Prähofer-Spohn) scaling function (black line) for comparison and the heat mode (right panel) with the $5/3$ Levy stable mode of maximal asymmetry (2.39) (black line), predicted by mode coupling equations for both $K=3$ and $K=4$ for comparison. Parameters for $K=3$: $\rho=1/4$, $a=4/3$, $E_{s}=1.16$ and for $K=4$: $\rho=1/4$, $a=1$, $c_{s}=7/4$, $E_{s}=1.11$, $c_{u}=0.8125$, satisfying the condition (2.37) for the absence of self-coupling. Two standard deviation neighborhoods are of the order of the line widths and are not shown for clarity. Other simulation parameters as in Fig. 4.
Figure 9: Comparison of umbiling scaling functions $f_{u}^{[K]}$ at different degeneracies $K$ (colored lines) with the best-fit KPZ (Prähofer-Spohn) scaling function (black line). We observe rapid convergence towards the latter with increasing degeneracy $K$. Scaling functions estimated from rescaled dynamical strucure factors as $f_{u}^{[K]}(r)\approx(E_{u}t_{K})^{2/3}S_{u}(c_{u}t_{K}+r(E_{u}t_{K})^{2/3},t_{K})$ at times $t_{2}=16000$ and $t_{3}=t_{4}=5000$. Inset shows difference of scaling functions heights (values at $r=0$) relative to the KPZ scaling function ($K=1$) as a function of the umbilic degeneracy $K$. Simulation parameters as in Fig. 8.
@@
Up to now we only described our numerical results for the double-degenerate $K=2$ umbilic mode, while our effective MCT theory predictions, namely Eqs. (2.36), (2.38), and (2.39) are expected to be valid for an umbilic mode of arbitrary degeneracy $K$. To test this prediction we have performed a similar analysis for $K=3$ and $K=4$ (see Fig. 8) which is more challenging numerically, when it comes to Monte-Carlo simulations, but has better finite-time convergence properties. The predictions for the dynamical exponents $z_{s}=5/3$ (2.38) and the scaling function (2.39) of the heat mode are found to be in excellent agreement with our numerical simulations, see right panel of Fig. 8. While the $z_{u}=3/2$ prediction of MCT (2.36) remains robust, the umbilic scaling function $f_{u}$ changes appreciably with the umbilic mode degeneracy $K$ as shown in the left panel of Fig. 8, cf. umbilic scaling function for $K=2$ in Fig. 4. With increasing $K$ the umbilic scaling functions $f_{u}^{[K]}$ starts to closely approach the KPZ scaling function, see Fig. 9. It would be interesting to investigate if the convergence $f_{u}^{[K]}\rightarrow f_{\rm KPZ}=f_{u}^{[1]}$ takes place in the “meanfield” limit $K\rightarrow\infty$.
@@ Acknowledgements
Appendix D
Effective NLFH equations for an umbilic mode with arbitrary higher degeneracy ($K>2$)
Analogously to the previous section, we consider an umbilic mode with arbitrary degeneracy $K$, appearing on the umbilic line of equal densities $\rho_{1}=\rho_{2}=\ldots=\rho_{K}=\rho_{K+1}=\rho$. The purpose of this section is to argue that also for higher degeneracy $K>2$ of the umbilic mode, we expect a universal umbilic dynamical structure factor. Because of the $K$-degeneracy of the umbilic mode, the diagonalizing transformation $R$ has K1K-1 free parameters parameters.
$R=(\rho(1-\rho))^{-\frac{1}{2}}\left(\begin{array}[]{c}v_{1}\\ v_{2}\\ \ldots\\ v_{K+1}\end{array}\right),$ (D.5)
Introducing basic vectors $e_{j}$ with elements $(e_{j})_{k}=\delta_{j,k}$, for $j,k=1,2,\ldots K+1$, we choose the orthonormal basis $v_{j}$ in (D.5) as
$v_{K+1}=\frac{1}{\sqrt{K+1}}(1,1,\ldots,1)=\frac{1}{\sqrt{K+1}}\sum_{j=1}^{K+1}e_{j},$ (D.6)
$v_{n}=\frac{1}{\sqrt{n+n^{2}}}\left(-ne_{n+1}+\sum_{j=1}^{n}e_{j}\right),\quad n=1,2,\ldots K.$ (D.7)
Under the transformation $R$, the NLFH equations for the umbilic mode $n=1,2,\ldots K$ take the following form
$\partial_{t}\phi_{s}=-\partial_{x}\left(c_{u}\phi_{s}-(s-1)b(K,a,\rho)A_{s}\phi_{s}^{2}+b(K,a,\rho)A_{s}\sum_{d=1}^{s-1}\phi_{d}^{2}\right$.
$\left.+2b(K,a,\rho)\phi_{s}\sum_{j=s+1}^{K}A_{j}\phi_{j}+D_{s}\partial_{x}\phi_{s}+B_{s}\xi_{s}\right)+\mbox{transitory terms},$ (D.8)
$\quad s=1,2,\ldots K,$
where the transitory terms of type $\phi_{s}\phi_{K+1}$ become negligibly small due to space separation between the umbilic $\phi_{s},\ s\leq K$ and single mode $\phi_{K+1}$ at late times. The coefficients $A_{s}$ in (D.8) are constants
$A_{s}=\sqrt{\frac{6}{s(s+1)}},$ (D.9)
while $b(K,a,\rho)$ is a function of system parameters,
$b(K,a,\rho)=-\sqrt{\frac{\rho(1-\rho)}{6}}\left(1+a+(K-2)a\rho\right).$ (D.10)
A Galilean transformation $(x,t)\rightarrow(x-c_{u}t,t)$ removes the terms $c_{u}\partial_{x}\phi_{s}$ in (D.8). Then, after rescaling time as $t\rightarrow\tau=b(K,a,\rho)t$, the reduced NLFH system (D.8) finally becomes
$\partial_{\tau}\phi_{s}=-\partial_{x}\left(-(s-1)A_{s}\phi_{s}^{2}+A_{s}\sum_{d=1}^{s-1}\phi_{d}^{2}+2\phi_{s}\sum_{j=s+1}^{K}A_{j}\phi_{j}+\tilde{D}_{s}\partial_{x}\phi_{s}+\tilde{B}_{s}\xi_{s}\right)$ (D.11)
$s=1,2,\ldots K$
The system (D.11) contains no dependence on system parameters $a,\rho$ (dependence on degeneracy $K$ enters through the number of equations), so we expect that the two-point correlation function for fixed $K$ is universal.
For $K=2$ the system (D.11) reduces to Eqs. (C.15) and (C.16). For $K=3$ and $K=4$ the NLFH equations take the form
$\partial_{\tau}\phi_{1}=-\partial_{x}\left(2\phi_{1}\left(\phi_{2}+\frac{1}{\sqrt{2}}\phi_{3}\right)+D_{1}\partial_{x}\phi_{1}+B_{1}\xi_{1}\right)$
$\partial_{\tau}\phi_{2}=-\partial_{x}\left(\phi_{1}^{2}-\phi_{2}^{2}+\sqrt{2}\,\phi_{2}\phi_{3}+D_{2}\partial_{x}\phi_{2}+B_{2}\xi_{2}\right)$ (D.12)
$\partial_{\tau}\phi_{3}=-\partial_{x}\left(\frac{1}{\sqrt{2}}\left(\phi_{1}^{2}+\phi_{2}^{2}-2\phi_{3}^{2}\right)+D_{3}\partial_{x}\phi_{3}+B_{3}\xi_{3}\right)$
and
$\partial_{\tau}\phi_{1}=-\partial_{x}\left(2\phi_{1}\left(\phi_{2}+\frac{1}{\sqrt{2}}\phi_{3}+\sqrt{\frac{3}{10}}\phi_{4}\right)+D_{1}\partial_{x}\phi_{1}+B_{1}\xi_{1}\right)$
$\partial_{\tau}\phi_{2}=-\partial_{x}\left(\phi_{1}^{2}-\phi_{2}^{2}+2\phi_{2}\left(\frac{1}{\sqrt{2}}\phi_{3}+\sqrt{\frac{3}{10}}\phi_{4}\right)+D_{2}\partial_{x}\phi_{2}+B_{2}\xi_{2}\right)$ (D.13)
$\partial_{\tau}\phi_{3}=-\partial_{x}\left(\frac{1}{\sqrt{2}}\left(\phi_{1}^{2}+\phi_{2}^{2}-2\phi_{3}^{2}\right)+2\phi_{3}\left(\sqrt{\frac{3}{10}}\phi_{4}\right)+D_{3}\partial_{x}\phi_{3}+B_{3}\xi_{3}\right)$
$\partial_{\tau}\phi_{4}=-\partial_{x}\left(\sqrt{\frac{3}{10}}\left(\phi_{1}^{2}+\phi_{2}^{2}+\phi_{3}^{2}-3\phi_{4}^{2}\right)+D_{4}\partial_{x}\phi_{4}+B_{4}\xi_{4}\right).$